# Fix-Con: Automatic Fault Localization and Repair of Deep Learning Model Conversions between Frameworks

### v3

@@ 7 Keras-to-Torch 36% 0% 89% 0%
VII Conclusion
We presented Fix-Con  an automated approach for fault localization and repair during model conversion between deep learning frameworks. Fix-Con is capable of detecting and fixing six commonly encountered fault types that are input and layer-based related to pre-processing, input dimensions, layer weights, biases, hyperparameters, and computation graph. We evaluated the effectiveness of Fix-Con in fixing model conversion bugs reported by Louloudakis et al. [63] for three widely used image recognition models converted across four different deep learning frameworks. We found that Fix-Con was highly effective in repairing the erroneous model conversions reported by completely fixing or significantly improving $1414$ out of the $1515$ cases under test, repairing a total of $462462$ bugs.

### v4

@@ 7 Keras-to-Torch 36% 0% 89% 0%
7. Challenges & Limitations
Upon FetaFix development, we encountered a number of challenges and limitation, which we describe in this section.
Node Matching Across Source & Target In order to perform the fault localization and repair process, FetaFix had to correctly match the nodes across Source and Target. However, the node matching was not one-to-one, due to (1) changes in graph structure applied by converter tools, (2) modifications in shape and internal structure of model nodes and (3) alterations in layer names between the model variants. To mitigate this problem, we introduced a solid set of requirements that had to be fulfilled for two nodes to be marked as the same. In detail, we required that two model nodes were (1) of the same type, (2) of the same shape, and (3) they shared a number of parameters on which values could only deviate by a small margin of difference (absolute value of $0.20.2$, obtained following step by step decrease to the point where the system failed to match the same nodes). For the moment however, we match only core computational layers (e.g., convolutions), considering that their number is guaranteed to remain unaffected by the conversion process due to model design. We also limit the matching algorithm to nodes containing weights and/or biases, as the presence of such properties increased significantly the matching ability of FetaFix, to the extent of more than 95% across Source and Target. We aim to improve our system to this direction by performing matching considering the presence of the same hyperparameters across nodes between Source and Target.
ONNX Opset Representation Limitations: For case #$1212$, we were unable to compile both Source and Target to the same opset, despite our efforts, due to ONNX operation incompatibilities. For that reason, we generated Source in opset=11 and Target in opset=13. However, the two opsets had completely different graph representations, and FetaFix was unable to correctly match the layers across model variants and perform fault localization and repair.
Computationally Intensive Fault Analysis: As the fault analysis we conducted performed per-element comparison for a number of inputs across each layer, it resulted in a vast amount of computations. This limited our tool from performing it to the full validation dataset, for reasons of tractability. For that matter, we selected a random sample of up to 100 images and we conducted our analysis on it, and repeated this process 3 times to solidify our results. We also accelerated the process by enabling parallel computation for the analysis, using Joblib python library (job, 2024).
8. Conclusion
We presented FetaFix, an automated approach for fault localization and repair during model conversion between deep learning frameworks. FetaFix is capable of detecting and fixing six commonly encountered fault types that are input and layer-based related to pre-processing, input dimensions, layer weights, biases, hyperparameters, and computation graph. We evaluated the effectiveness of FetaFix in fixing previously reported model conversion bugs (Louloudakis et al., 2023a) for three widely used image recognition models converted across four different deep learning frameworks. We found that FetaFix was highly effective in repairing the erroneous model conversions reported by completely fixing or significantly improving $1414$ out of the $1515$ cases under test, repairing a total of $462462$ bugs.
Acknowledgements.
The research has been funded by the Royal Society Industry Fellowship ”AutoTest: Testing Autonomous Vehicle Perception Safety on Hardware Accelerators” and Huawei Edinburgh Joint Lab Project ”RobustCheck: Testing Robustness of Compiler Optimisations and Deep Learning Frameworks”.